OpenStudy (anonymous):

solve the silmutaneous question 3x+y=2 and 3x^2 +y^2+xy=6

OpenStudy (anonymous):

ick solve the first one for y and then substitute in the second one

OpenStudy (anonymous):

\[y=2-3x\] \[3x^2 +y^2+xy=6 \] \[3x^2+(2-3x)^2+x(2-3x)=6\]

OpenStudy (anonymous):

bunch of algebra gives the left hand side as \[9x^2-10x+4\] so you have \[9x^2-10x+4=6\] or \[9x^2-10x-2=0\] quadratic equation is all that is left to do